OpenStudy (anonymous):

Find the value of three times the quantity x minus two when x equals eight. 1 eighteen twenty-two twenty-four

OpenStudy (anonymous):

3x-2 when x=8 3*8-2 =24-2 =22 twenty-two :)

OpenStudy (anonymous):

thankz datz wat i wus thinking